# On the future cover of a sofic shift

### v2

@@ 4.2. Subset constructions for sofic shifts
One of the possible applications of Theorem 4.6 is to the minimal right-resolving presentations of two conjugate irreducible sofic shifts. The paper finishes with three examples where the canonical cover $(G^{\prime},L_{G^{\prime}})$ arising in this setup is exhibited and compared to the future cover. For this we need a general procedure to construct the future cover.
5. Constructing the future cover
In this section we present a method to construct the future cover of a sofic shift from any given presentation of it.
5.1. On the merging of a labeled graph
Let $(H,L_{H})$ be a labeled graph. The merged graph $([H],L_{[H]})$ of $(H,L_{H})$ is then defined as follows, cf. page 78 of [LM]. The set of vertexes $V_{[H]}$ is the set
$V_{[H]}:=\left\{f_{H}(v):\ v\in V_{H}\right\}$
and for $a\in A$ there is a labeled arrow $f_{H}(v)\overset{a}{\to}f_{H}(w)$ when there is are vertexes $v^{\prime},w^{\prime}\in V_{H}$ such that $f_{H}(v^{\prime})=f_{H}(v),\ f_{H}(w^{\prime})=f_{H}(w)$ and there is a labeled arrow $v^{\prime}\overset{a}{\to}w^{\prime}$ in $H$. When $(H,L_{H})$ is right-resolving we can define a map $f_{H}:E_{H}\to E_{[H]}$ such that $f_{H}(e)$ is the arrow $f_{H}(e)\in E_{[H]}$ with the properties
  • $s_{[H]}(f_{H}(e))=f_{H}\left(s_{H}(e)\right)$, and
  • $L_{[H]}(e)=L_{H}(e)$.
Then $f_{H}:H\to[H]$ is a labeled-graphs homomorphism.
Lemma 5.1.
Assume that $(H,L_{H})$ is right-resolving. Let $v,w\in V_{H}$. There is a labeled arrow $f_{H}(v)\overset{a}{\to}f_{H}(w)$ in $[H]$ if and only if there is a labeled arrow $v\overset{a}{\to}u$ in $H$ such that $f_{H}(u)=f_{H}(w)$.
Proof.
Assume that $f_{H}(v)\overset{a}{\to}f_{H}(w)$ in $[H]$ and let $v_{1},w_{1}\in V_{H}$ be such that $v_{1}\overset{a}{\to}w_{1}$ in $H$, $f_{H}(v_{1})=f_{H}(v)$ and $f_{H}(w_{1})=f_{H}(w)$. Since $f_{H}(v)=f_{H}(v_{1})$ there is an arrow $e$ in $E_{H}$ labeled $a$ which leaves $v$. Set $u:=t_{H}(e)$. By using that $(H,L_{H})$ is right-resolving we find then that
$f_{H}(u)=\frac{f_{H}(v)}{a}=\frac{f_{H}(v_{1})}{a}=f_{H}(w_{1}).$
This establishes the ’only if’ part. The ’if’ part is trivial. ∎
Lemma 5.2.
Assume that $(H,L_{H})$ is right-resolving. Then $([H],L_{[H]})$ is right-resolving and follower-separated, and there is a labeled-graph homomorphism $f_{H}:H\to[H]$ such that $f_{H}:X_{H}\to X_{[H]}$ is a factor map.
Proof.
$([H],L_{[H]})$ is right-resolving by Lemma 3.3.8 of [LM]. To see that $f_{H}:X_{H}\to X_{[H]}$ is surjective, consider a finite path
(17) $f_{H}(v_{1})\overset{a_{1}}{\to}f_{H}(v_{2})\overset{a_{2}}{\to}f_{H}(v_{3})\overset{a_{3}}{\to}\cdots\cdots\overset{a_{n-1}}{\to}f_{H}(v_{n})$
in $[H]$. Repeated applications of Lemma 5.1 give a path
$v_{1}\overset{a_{1}}{\to}u_{2}\overset{a_{2}}{\to}u_{3}\overset{a_{3}}{\to}\cdots\cdots\overset{a_{n-1}}{\to}u_{n}$
in $H$ which is mapped to (17) under $f_{H}$. The surjectivity of $f_{H}$ follows therefore from compactness of $X_{H}$. The argument shows also that the follower set of $f_{H}(v_{1})$ in $([H],L_{[H]})$ is the same as the follower set of $v_{1}$ in $(H,L_{H})$; in symbols
$f_{[H]}(f_{H}(v_{1}))=f_{H}(v_{1}).$
In particular, $([H],L_{[H]})$ is follower-separated ∎
We note that the morphism $f_{H}:H\to[H]$ is injective on $V_{H}$ if and only if it is an isomorphism of labeled graphs, and this happens if and only if $(H,L_{H})$ is follower-separated.
5.2. A cover of the future cover
Let $(G,L_{G})$ be a labeled graph and $Y=L_{G}(X_{G})$ the sofic shift presented by $(G,L_{G})$. When $D\subseteq V_{G}$ and $a\in A$, set
$D,a]:=\left\{t_{G}(e):\ e\in E_{G},\ s_{G}(e)\in D,\ L_{G}(e)=a\right\};$
the set of vertexes that can be reached from an element of $D$ using an edge in $G$ labeled by $a$. We consider the collection $2^{V_{G}}$ of non-empty subsets of $V_{G}$ as the set of vertexes in a graph $\overline{G}$ with edges
$E_{\overline{G}}:=\left\{(D,a)\in 2^{V_{G}}\times A:\ [D,a]\neq\emptyset\right\}.$
The start vertex $s_{\overline{G}}(D,a)$ of $(D,a)\in E_{\overline{G}}$ is $D\in 2^{V_{G}}$ and the terminal vertex $t_{\overline{G}}(D,a)$ is $[D,a]$. We define a labeling $L_{\overline{G}}$ of $\overline{G}$ such that
$L_{\overline{G}}(D,a):=a$
when $(D,a)\in E_{\overline{G}}$. Then $(\overline{G},L_{\overline{G}})$ is a right-resolving labeled graph and $L_{\overline{G}}(X_{\overline{G}})=Y$.
For $y\in Y$, set
$D^{y}:=\left\{t_{G}(x):\ x\in X_{G}(-\infty,-1],\ L_{G}(x)=y_{(-\infty,-1]}\right\}.$
The set of vertexes $V_{\underline{G}}:=\left\{D^{y}:\ y\in Y\right\}$ is a hereditary subset of vertexes in $V_{\overline{G}}$; that is, if $e\in E_{\overline{G}}$ and $s_{\overline{G}}(e)\in V_{\underline{G}}$, then $s_{\overline{G}}(e)\in V_{\underline{G}}$. Hence, when we set $E_{\underline{G}}:=\left\{e\in E_{\overline{G}}:\ s_{\overline{G}}(e)\in V_{\underline{G}}\right\}$, we have defined a graph $\underline{G}$, and hence also a labeled graph $(\underline{G},L_{\underline{G}})$, when $L_{\underline{G}}:=L_{\overline{G}}|_{E_{\underline{G}}}$.
Note that we can define a shift-commuting map $\alpha_{G}:Y\to X_{\underline{G}}$ such that
  • a)
    $s_{\underline{G}}(\alpha_{G}(y)_{k})=D^{\sigma^{k}(y)}$ for all $k\in\mathbb{Z}$, and
  • b)
    $L_{\underline{G}}\left(\alpha_{G}(y)\right)=y$.
In particular, $L_{\underline{G}}(X_{\underline{G}})=Y$ and $(\underline{G},L_{\underline{G}})$ is thus a presentation of $Y$. We show next that the future cover of $Y$ is the merged graph of $(\underline{G},L_{\overline{G}})$.
Let $([\underline{G}],L_{[\underline{G}]})$ be the merged graph of $(\underline{G},L_{\underline{G}})$. Thus the set of vertexes $V_{[\underline{G}]}$ consists of the set
$V_{[\underline{G}]}:=\left\{f_{\underline{G}}(v):\ v\in V_{\underline{G}}\right\}.$
By Lemma 5.2 $([\underline{G}],L_{[\underline{G}]})$ is right-resolving and the morphism $f_{\underline{G}}:\underline{G}\to[\underline{G}]$ of labeled graphs induces a factor map $f_{\underline{G}}:X_{\underline{G}}\to X_{[\underline{G}]}$.
Lemma 5.3.
$F(y)=f_{\underline{G}}(D^{y})$ for all $y\in Y$.
Proof.
It follows from the definitions that
$f_{\underline{G}}(D^{y})=\left\{L_{\underline{G}}(x):\ x\in X_{\underline{G}}[0,\infty),\ s_{\underline{G}}(x)=D^{y}\right\}$
$=\left\{L_{G}(x^{\prime}):\ x^{\prime}\in X_{G}[0,\infty),\ s_{G}(x^{\prime})\in D^{y}\right\}$
$=\left\{z_{[0,\infty)}:\ z\in Y,\ z_{j}=y_{j},\ j\leq-1\right\}=F(y).$
Lemma 5.4.
Let $y,z\in Y,\ a\in A$. There is a labeled arrow
(18) $F(y)\overset{a}{\to}F(z)$
in $\mathbb{K}(Y)$ if and only if there is a labeled arrow
(19) $f_{\underline{G}}(D^{y})\overset{a}{\to}f_{\underline{G}}(D^{z})$
in $\underline{G}]$.
Proof.
By Lemma 5.3 an arrow as in (18) exists if and only if $f_{\underline{G}}(D^{z})=\frac{f_{\underline{G}}(D^{y})}{a}$, or if and only if
$\left\{L_{\underline{G}}(x):\ x\in X_{\underline{G}}[0,\infty),\ s_{\underline{G}}(x)=D^{z}\right\}=\frac{\left\{L_{\underline{G}}(x^{\prime}):\ x^{\prime}\in X_{\underline{G}}[0,\infty),\ s_{\underline{G}}(x^{\prime})=D^{y}\right\}}{a}.$
The latter equality holds exactly when there is a labeled arrow $D^{y}\overset{a}{\to}D$ in $\underline{G}$ for some $D\in V_{\underline{G}}$ with $f_{\underline{G}}(D)=f_{\underline{G}}(D^{z})$. By Lemma 5.1 this happens if and only there is a labeled arrow as in (19). ∎
It follows from the last two lemmas that there is an isomorphism of labeled graphs $(\mathbb{K}(Y),L_{\mathbb{K}(Y)})\simeq([\underline{G}],L_{[\underline{G}]})$ defined such that
$V_{\mathbb{K}(Y)}\ni F(y)\mapsto f_{\underline{G}}(D^{y}),$
allowing us to make the following identification:
Corollary 5.5.
$(\mathbb{K}(Y),L_{\mathbb{K}(Y)})=([\underline{G}],L_{[\underline{G}]})$ as labeled graphs.
6. Three examples
It seems to be a well-established tradition that examples of sofic shifts start with the even shift, and we have no reason to deviate from this. The minimal right-resolving presentation of the even shift is shown in Figure 5 and the future cover in Figure 3. When $(G,L_{G})$ is the labeled graph from Figure 5, the corresponding labeled graph $(G^{\prime},L_{G^{\prime}})$ from Theorem 4.6 is
@@ 6. Three examples
If instead $(G,L_{G})$ is the labeled graph from Figure 3, the labeled graph $(G^{\prime},L_{G^{\prime}})$ becomes
$\{a,b,c\}$$\{a,b\}$$\{a\}$$\{b\}$$\{a,c\}$000111100
Thus we see, already from the even shift, that the minimal right-resolving cover, the future cover, the labeled graph $(G^{\prime},L_{G^{\prime}})$ from Theorem 4.6 applied first to the minimal right-resolving cover and then to the future cover, are all different covers.
We end the paper by giving two examples where the future cover of a mixing sofic shift is the same as the minimal right-resolving cover, while the canonical cover from Theorem 4.6 applied to the minimal right-resolving cover is different. The examples also show that in some cases the labeled graph $(\underline{G},L_{\underline{G}})$ from Section 5.2 is the same as $(G^{\prime},L_{G^{\prime}})$ and in others it isn’t.
Example 6.1.
The following labeled graph $(G,L_{G})$ is irreducible, right-resolving and follower-separated and hence the minimal right resolving graph of the sofic shift $Y$ it represents. Note that $Y$ is mixing.
(20) aabb112201122
The graph $(\underline{G},L_{\underline{G}})$ is the following.
$\{a\}$$\{b\}$$\{a,b\}$11220112222110
Here the vertex $\{a,b\}$ has the same follower set as $\{a\}$ and the merged graph, which is the future cover $(\mathbb{K}(Y),L_{\mathbb{K}(Y)})$ by Corollary 5.5, is the graph (20) we started with. The future cover is therefore the same as the minimal right-resolving cover. In this case the labeled graph $G^{\prime}$ is
$\{a\}$$\{b\}$$\{a,b\}$1122011222211
Thus $(G^{\prime},L_{G^{\prime}})$ is a proper subgraph of $(\underline{G},L_{\underline{G}})$ in this case.
Example 6.2.
The following labeled graph $(G,L_{G})$ is irreducible, right resolving and follower-separated and hence the minimal right resolving graph of the sofic shift $Y$ it represents. Note that $Y$ is mixing.
(21) aabb0220111122
In this case the graphs $(\underline{G},L_{\underline{G}})$ and $(G^{\prime},L_{G^{\prime}})$ are the same:
$\{a\}$$\{b\}$$\{a,b\}$022011112202211
Since the vertex $\{a,b\}$ has the same follower set as $\{a\}$, the future cover is again the labeled graph (21) we started with, and hence equal to the minimal right-resolving cover, while the canonical cover $(G^{\prime},L_{G^{\prime}})$ is bigger.

### v3

@@ 4.2. Subset constructions for sofic shifts
For a sofic shift $Y$, set
$(\mathbb{K}^{\prime}(Y),L_{\mathbb{K}^{\prime}(Y)}):=(\mathbb{K}(Y)^{\prime},L_{\mathbb{K}(Y)^{\prime}}).$
By Theorem 2.11 and Theorem 4.6 this defines a strongly canonical cover for sofic shifts:
Theorem 4.7.
Let $\psi:Y\to Z$ be a conjugacy of sofic shifts. There is a unique conjugacy $\psi_{\mathbb{K}^{\prime}}:X_{\mathbb{K}^{\prime}(Y)}\to X_{\mathbb{K}^{\prime}(Z)}$ such that
${X_{\mathbb{K}^{\prime}(Y)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi_{\mathbb{K}^{\prime}}}$$\scriptstyle{L_{\mathbb{K}^{\prime}(Y)}}$${X_{\mathbb{K}^{\prime}(Z)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{L_{\mathbb{K}^{\prime}(Z)}}$${Y\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi}$${Z}$
commutes.
For an irreducible sofic shift $Y$ we set
$(\mathbb{F}^{\prime}(Y),L_{\mathbb{F}^{\prime}(Y)}):=(\mathbb{F}(Y)^{\prime},L_{\mathbb{F}(Y)^{\prime}}).$
By Corollary 10 and Theorem 4.6 this defines also a strongly canonical cover for irreducible sofic shifts:
Theorem 4.8.
Let $\psi:Y\to Z$ be a conjugacy of irreducible sofic shifts. There is a unique conjugacy $\psi_{\mathbb{F}^{\prime}}:X_{\mathbb{F}^{\prime}(Y)}\to X_{\mathbb{F}^{\prime}(Z)}$ such that
${X_{\mathbb{F}^{\prime}(Y)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi_{\mathbb{F}^{\prime}}}$$\scriptstyle{L_{\mathbb{F}^{\prime}(Y)}}$${X_{\mathbb{F}^{\prime}(Z)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{L_{\mathbb{F}^{\prime}(Z)}}$${Y\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi}$${Z}$
commutes.
Example 4.9.
It seems to be a well-established tradition that examples of sofic shifts start with the even shift, and we have no reason to deviate from this. The minimal right-resolving presentation $(\mathbb{F}(E),L_{\mathbb{F}(E)})$ of the even shift $E$ is shown in Figure 5 and the future cover $(\mathbb{K}(E),L_{\mathbb{K}(E)})$ in Figure 3. Then $(\mathbb{F}^{\prime}(E),L_{\mathbb{F}^{\prime}(E)})$ is the following labeled graph,
@@
and $(\mathbb{K}^{\prime}(E),L_{\mathbb{K}^{\prime}(E)})$ is the labeled graph
$\{a,b,c\}$$\{a,c\}$$\{a\}$$\{b\}$$\{b,c\}$000111100
We see therefore, already from the even shift, that in general the canonical covers $(\mathbb{F}(E),L_{\mathbb{F}(E)})$, $(\mathbb{K}(E),L_{\mathbb{K}(E)})$, $(\mathbb{F}^{\prime}(E),L_{\mathbb{F}^{\prime}(E)})$ and $(\mathbb{K}^{\prime}(E),L_{\mathbb{K}^{\prime}(E)})$ are all different.